# Pressure2Motion: Hierarchical Motion Synthesis from Ground Pressure with Text Guidance

### v1

@@ 5 Additional Ablation Study
The results of this experiment are shown in Table 3, where we compare the full model with versions that exclude ControlNet and Adapter. For the version without ControlNet, we observe a significant increase in FID (1.3683) and MPJPE (0.1951), indicating that removing the ControlNet impairs the model’s ability to properly align the motion with the pressure signals, resulting in less accurate motion generation. Similarly, removing the Adapter leads to a noticeable degradation in performance, with FID increasing to 0.695 and MPJPE rising to 0.2092. These results demonstrate the critical role of both components in ensuring the high fidelity and physical plausibility of the generated motions. Moreover, concatenating the two pressure features (Movement Trajectories and Posture Shifts) directly and feeding them into either the ControlNet or Adapter results in inferior performance compared to the full model. This suggests that the hierarchical structure, where ControlNet handles the overall movement trajectory and the Adapter fine-tunes the posture shifts, is essential for generating realistic and semantically aligned motions.
Figure 9 provides visual comparisons of the motion sequences generated by the different model variations. These visualizations further confirm that the full model consistently produces motions that are more physically plausible and aligned with the pressure data, especially in areas such as foot-ground contact.
@@ 6 Limitations
Another major limitation of our model lies in its computational complexity. While our approach demonstrates high fidelity in synthesizing pressure-aware motions, the underlying architecture—specifically the pressure feature extraction module and the hierarchical pressure-modulated motion synthesis framework—is relatively large and computationally demanding. Additionally, the motion generation process, based on the diffusion model, involves multiple denoising steps, making the inference process slower. On a single NVIDIA A800 GPU, generating one motion sequence takes approximately 180 seconds. This extended inference time can become a bottleneck when deploying the model in real-time applications. As a future direction, we plan to investigate strategies, such as more efficient pressure feature extraction techniques, and explore inference optimization to speed up the motion generation process while maintaining high-quality results.

### v2

@@ 10 Additional Ablation Study
The results of this experiment are shown in Table 3, where we compare the full model with versions that exclude ControlNet and Adapter. For the version without ControlNet, we observe a significant increase in FID (1.3683) and MPJPE (0.1951), indicating that removing the ControlNet impairs the model’s ability to properly align the motion with the pressure signals, resulting in less accurate motion reconstruction. Similarly, removing the Adapter leads to a noticeable degradation in performance, with FID increasing to 0.695 and MPJPE rising to 0.2092. These results demonstrate the critical role of both components in ensuring the high fidelity and physical plausibility of the reconstructed motions. Moreover, concatenating the two pressure features (Movement Trajectories and Posture Shifts) directly and feeding them into either the ControlNet or Adapter results in inferior performance compared to the full model. This suggests that the hierarchical structure, where ControlNet handles the overall movement trajectory and the Adapter fine-tunes the posture shifts, is essential for reconstructing realistic and semantically aligned motions.
Figure 9 provides visual comparisons of the motion sequences reconstructed by the different model variations. These visualizations further confirm that the full model consistently produces motions that are more physically plausible and aligned with the pressure data, especially in areas such as foot-ground contact.
Table 3: Additional Ablation study of Model Archtecture.
[…]
w/o ctrlnet 1.3683 0.0621 0.6120 0.1694 0.1951
w/o Adapter 0.695 0.0634 0.6655 0.1702 0.2092
Full 0.262 0.0553 0.4260 0.1273 0.1622
11 More Visualization Results
We present more motion reconstruction results to further demonstrate the effectiveness of our method in reconstructing human motion from sparse pressure data. Figure 10 showcases a variety of reconstructed motions across different scenarios, highlighting the robustness of our approach in the pressure-to-motion task. The first few cases demonstrate typical human actions, such as walking, standing, and some daily activities. Our model reconstructs high-fidelity motions that align well with the pressure data and text prompts, maintaining both physical consistency and semantic plausibility.
Particularly interesting are the last two cases in Figure 10. The second-to-last case corresponds to a jumping motion, where no pressure is applied during the jump. Despite the absence of pressure in the air, our model successfully reconstructs a realistic jumping motion, demonstrating its ability to handle scenarios with no foot-ground contact.
The final case in Figure 10 shows a dynamic plank position, which involves complex pressure distributions from both the hands and feet. Our method effectively handles this rare and specialized pressure contact scenario, reconstructing a physically plausible motion that corresponds to the simultaneous pressure from all four limbs. This demonstrates the versatility of our model in handling uncommon or intricate pressure patterns.
@@ 12 Limitations
Another major limitation of our model lies in its computational complexity. While our approach demonstrates high fidelity in reconstructing pressure-aware motions, the underlying architecture—specifically the pressure feature extraction module and the hierarchical pressure-modulated motion reconstruction framework—is relatively large and computationally demanding. Additionally, the motion reconstruction process, based on the diffusion model, involves multiple denoising steps, making the inference process slower. On a single NVIDIA A800 GPU, reconstructing one motion sequence takes approximately 180 seconds. This extended inference time can become a bottleneck when deploying the model in real-time applications. As a future direction, we plan to investigate strategies, such as more efficient pressure feature extraction techniques, and explore inference optimization to speed up the motion reconstruction process while maintaining high-quality results.
A third limitation concerns the nature of our textual guidance. The MPL dataset’s text descriptions are entirely generated by a VLM. While this provides consistent and detailed annotations at scale, VLM-generated text tends to be homogeneous and overly descriptive, adhering to a specific stylistic pattern (e.Sg., ”The person raises their left arm”). This clean, literal data distribution does not reflect the full diversity and ambiguity of real-world human language. Human prompts are often more abstract, colloquial, underspecified, or describe high-level goals rather than explicit kinematics (e.g., ”Look for something on the floor” vs. ”Bend over and turn head”). Consequently, our model may be overfitted to this VLM-specific text style and less robust to ”in-the-wild” human-authored prompts.
Future work will focus on addressing these limitations by: 1) Enhancing robustness and expanding the dataset to include more dynamic activities (e.g., inclined surfaces); 2) Investigating inference optimization strategies to enable real-time applications; 3) Improving robustness to diverse, human-authored text; and 4) Leveraging pressure data in simulated reinforcement learning or personalized motion modeling.